A certain radioactive element decays exponentially with a half-life of 5 hours. Find the time that has elapsed if a 50mg sample has decayed to 17 mg.
Answer by stanbon(75887)

A certain radioactive element decays exponentially with a half-life of 5 hours. Find the time that has elapsed if a 50mg sample has decayed to 17 mg.
--------------------
A(t) = A(0)(1/2)^(t/5)
17 = 50 (1/2)^(t/5)
0.34 = (1/2)^(t/5)
Take the log of both sides to get:
(t/5)log(1/2) = log(0.34)
t/5 = 1.5564
t = 7.782 hours
================
Cheers,
Stan H.